# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 8.2 Regularizing the graph and basic analysis
We will first ensure that $G$ has nice properties, before the “heavy-duty” branching rules. Specifically, we say $G$ is in 4-base position (or 4BP for short) if the following conditions hold:
@@
  • $G$ has no connected-components of size larger than some constant $K$.
Note: Let us explain the last clause in the definition of 4BP, and the choice of constant $K$ (which we will not specify explicitly). In light of Lemma 35, our branching algorithms will generate principal subproblems $G-(U_{0}\cup U_{1})$ where $\Delta k=|U_{1}|$ is lower-bounded by some constant value $c$. Such subproblems may be produced directly (e.g. by guessing a vertex is in the cover), or may be obtained by applying (P1) to a principal subproblem $G-X$. If the constant $K$ is large enough (in particular, larger than $8c$), then any such principal subproblem $G-(U_{0}\cup U_{1})$ with drop $(\Delta\mu,c)$ automatically has drop $(\Delta\mu,c^{*})$ “for free”. For example, this principle was used in Proposition 36; we will use it implicitly through the remainder of the algorithm analysis.
@@
If $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 2$, but $G$ is not in 4BP, then it has available branch-seq $[(0,2)]$, $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(1,3^{*}),(1,3^{*})]$
@@
If $G$ has a connected-component $X$ of constant size, we may solve vertex cover on it exactly in constant time, and update $k^{\prime}=k-|C_{X}|$ where $C_{X}$ is a good cover of $X$. Since $\operatorname{minsurp}(G)\geq 2$, we have $|C_{X}|\geq 2$, and so removing this component gives drop $[(0,2)]$.
If $G$ has a funnel $u$ with out-neighbor $x$, then we branch on subproblems $G-N[u],G-N[x]$; these both have drop $(1,3^{*})$ by Lemma 35. Since $\operatorname{minsurp}(G)\geq 2$, rules (P1) and (P2) cannot be applied; so for the remainder of the analysis we may suppose that $G$ is simplified.
@@
Finally, if $G$ has a 4-cycle $u_{1},u_{2},u_{3},u_{4}$, we apply (A3) to get subproblems $G-\{u_{1},u_{3}\}$ and $G-N(u_{1},u_{3})$. We claim these subproblems both have drops $(1,3^{*})$ after simplification. Let us first show this holds for subproblem $G^{\prime}=G-\{u_{1},u_{3}\}$, which has drop $(1,2)$ directly. If $u_{2}\not\sim u_{4}$, then Observation 16 immediately gives $S(G^{\prime})\geq 2$. So suppose $u_{2}\sim u_{4}$; since $G$ is simplified, this is only possible if $u_{1},u_{2},u_{3},u_{4}$ all have degree 4 (else they would have a 3-triangle). Then $u_{1}\not\sim u_{3}$ as otherwise $u_{1}$ would have a funnel with its neighbors $u_{2},u_{3},u_{4}$. So $G^{\prime}$ has at least 4 subcubic vertices (the neighbors of $u_{1}$) and it has a 2-vertex $u_{2}$, so Proposition 33 with $s=1$ gives $S(G^{\prime})\geq 1^{*}$.
We next argue that $G^{\prime\prime}=G-N(u_{1},u_{3})$ has drop $(1,3^{*})$. If $\operatorname{codeg}(u_{1},u_{3})\geq 3$, then $G^{\prime\prime}$ has drop $(1,3^{*})$ directly by Lemma 35. Otherwise, if $\operatorname{codeg}(u_{1},u_{3})=2$, then $G^{\prime\prime}=G-\{u_{2},u_{4}\}$; this has drop $(1,3^{*})$ by precisely the same reasoning as for the graph $G^{\prime}=G-\{u_{1},u_{3}\}$. ∎
@@
If $G$ is in 4BP and $u$ is a 4-vertex, then $G-N[u]$ has at least 8 subcubic vertices.
@@
Suppose $u$ has 3-neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{s}$ We have $\operatorname{codeg}(u,v_{i})=0$ for $i\leq s$, since otherwise $v_{i}$ would have a 3-triangle. Likewise, we have $\operatorname{codeg}(u,v_{i})\leq 1$ for $i>s$, as otherwise $G$ would have a 4-cycle. Thus, the neighborhood of $u$ has at most $\lfloor(4-s)/2\rfloor$ edges, and hence at least $3s+4(4-s)-2\lfloor(4-s)/2\rfloor-4\geq 8$ edges from $N(u)$ to vertices $t$ at distance two from $u$. Since $G$ has no 4-cycles, each such vertex $t$ has an edge to precisely one vertex in $N(u)$. ∎
We now analyze the easier steps of the algorithm line-by-line. Note that, starting at Line 4, the graph $G$ is 4-regular. Lines 3 and 5, which are more involved, will be considered later.
@@
First suppose that $\operatorname{bonus}(G)=0$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\geq 5$, then we apply Section 5; otherwise, we apply Proposition 37. We need:
@@
Finally suppose $\operatorname{bonus}(G)=\alpha$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\leq 4$, then we can apply Proposition 37. Finally, suppose $G$ is simplified and has a vertex $u$ of degree at least $5$. If $\operatorname{shad}(N[u])\leq 4-\deg(u)$, then we apply Lemma 26. If $\operatorname{shad}(N[u])\geq 5-\deg(u)$, then we split on $u$; subproblem $G-N[u]$ has drop $(2,5)$, while subproblem $G-u$ retains the subcubic vertex in the original graph and so has drop $(0.5,1^{*})$. Over all cases, we need:
@@
Line 4. Let $N(u)=\{v_{1},\dots,v_{4}\}$ where $v_{1}\sim v_{2}$. Since $G$ has no 4-cycles, $v_{1}\not\sim\{v_{3},v_{4}\}$. We split on $v_{3}$, getting subproblems $G-N[v_{3}],G-v_{3}$ with drops $(1.5,4^{*})$ and $(0.5,1^{*})$. In the latter subproblem, we apply Proposition 30 to the funnel $u$ with out-neighbor $v_{4}$ and vertex $v_{1}\in N(u)\setminus N(v_{4})$, where $r=\deg_{G-v_{3}}(v_{1})+\deg_{G-v_{3}}(v_{4})-\operatorname{codeg}_{G-v_{3}}(v_{1},v_{4})-1\geq 4+3-1-1=5$. We need:
@@ 8.3 Line 3: Branching on a 4-vertex with a 3-neighbor
Suppose Line 3 selects a 4-vertex $u$, with 3-neighbors $v_{1},\dots,v_{s}$ and $4$-neighbors $v_{s+1},\dots,v_{4}$. Let $x_{1},x_{2}$ be the other two neighbors of $v_{1}$, sorted so $\deg(x_{1})\geq\deg(x_{2})$. Since $G$ has no 4-cycles, and the 3-vertices $v_{1},\dots,v_{s}$ have no triangles, all vertices $u,v_{i},x_{i}$ are distinct, and $\{x_{1},x_{2}\}\not\sim\{u,v_{2},v_{3},v_{4}\}$. Note that there may be multiple ways of ordering vertices in this way; e.g., we can permute the vertices $v_{1},\dots,v_{s}$ or $v_{s+1},\dots,v_{4}$, or if if $\deg(x_{1})=\deg(x_{2})$ we can swap $x_{1}/x_{2}$.
Let us fix some ordering and consider the graphs $G_{0}=G-N[u]$ and $G_{1}=G-u$ produced by splitting on $u$. There are two strategies we can use for further branching. First, we can apply (P2) to the graph $G_{1}$, contracting its 2-vertices $v_{1},\dots,v_{s}$ into vertices $z_{1},\dots,z_{s}$; we let $G_{1}^{\prime}$ denote the resulting graph, and in this case we overall generate two subproblems $G_{0},G_{1}^{\prime}$. Second, we can apply (A1) to the 2-vertex $v_{1}$ in $G-u$, generating further subproblems $G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[x_{1},x_{2}]$; so in this case we overall generate three subproblems $G_{0},G_{2},G_{3}$.
Throughout, we write $r=\deg(x_{1})+\deg(x_{2})-2=\deg_{G_{1}^{\prime}}(z_{1})$.
@@
For the first result, if $s\leq 3$, then $G-N[v_{1}]$ has a 3-vertex $v_{4}$ and a 2-vertex $v_{2}\not\sim v_{4}$; applying (P2) to $v_{2}$ reduces $k$ by one, and still leaves a subcubic vertex $v_{4}$. If $s=4$, then $G-N[v_{1}]$ has non-adjacent 2-vertices $v_{2},v_{3}$, so $S(G_{2})\geq 2$ by Observation 11.
For the next two results, recall from Proposition 38 that $G_{0}$ has at least 8 subcubic vertices. If $r=4$, then $\deg_{G_{0}}(x_{1})=\deg_{G_{0}}(x_{2})=2$ and $x_{1}\not\sim x_{2}$, so by Proposition 34 we have $S(G_{0})\geq 2^{*}$. Similarly, if $r=5$, then $\deg_{G_{0}}(x_{2})=2$ so by Proposition 33 we have $S(G_{0})\geq 1^{*}$. ∎
@@
Suppose $r\geq 5$. Then $G_{3}$ is a principal subproblem with excess two and $\Delta k=r+2$ and $\operatorname{minsurp}(G_{3})\geq 4-r$. Furthermore, the following bounds all hold:
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$, then $R(N[x_{1}])\geq 2^{*}$.
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$ and $s=1$, then $R(N[x_{1}])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-2$, then $R(N[x_{1}])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-1$, then $G-N[x_{1}]$ is $3^{*}$-vul.
@@
Let $H=G-u-N[x_{1},x_{2}]$. Since $G$ has no 4-cycles, we have $|N[x_{1},x_{2}]|=r+1$ and so $H$ is principal with excess two and $\Delta k=r+2$. Since $\deg(x_{1})\geq\deg(x_{2})$, we have $\deg(x_{1})=4$ and $\deg(x_{2})=r-2$. Note that if $s=1$, then the sole 3-neighbor of $u$ gets removed from $H$, and so $\operatorname{mindeg}(H)\geq 1$. Also note that, by Proposition 38, the graph $G-N[x_{1}]$ has at least 8 subcubic vertices, and since $G$ is in 4BP we have $\operatorname{shad}(N[x_{1}])\geq 1$.
Let $I$ be a min-set in $H$ and $\ell=\operatorname{surp}_{H}(I)$. If $\ell\leq-2$ or $s=1$ then necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq\ell+(\deg_{G}(x_{2})-1)=\ell+r-3$. Since $G$ is in 4BP, we have $\operatorname{shad}(N[x_{1}])\geq 1$ so $\ell\geq 4-r$. Furthermore, if $\ell=4-r$ exactly, then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})=1$, and then Proposition 33 applied to indset $I\cup\{x_{2}\}$ gives $R(N[x_{1}])\geq 2^{*}$ for $|I|=1$ and $R(N[x_{1}])\geq 3^{*}$ for $|I|\geq 2$.
Finally, suppose $r=6$ and $\ell=-1$. If $|I|=1$, then $H$ has an isolated vertex $t$; since $G$ has no 4-cycles, this is only possible if $\deg_{G}(t)=3$ and $t\sim u$ and $\operatorname{codeg}(t,x_{1})=\operatorname{codeg}(t,x_{2})=1$. In this case, $\deg_{G-N[x_{1}]}(t)=2$ and by Proposition 33 we have $R(N[x_{1}])\geq 1^{*}$. Otherwise, we have $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq 2$, which by Proposition 36 implies that $G-N[x_{1}]$ is $3^{*}$-vul. ∎
@@
Case I: $\boldsymbol{s\geq 2}$ and $\boldsymbol{S(G_{0})\geq 2^{*}}$. We generate subproblems $G_{0},G_{1}^{\prime}$. If $s\in\{2,3\}$, then $G_{1}^{\prime}$ has drop $(0.5,3)$ or $(0.5,4)$ and also has a 3-vertex $v_{4}$; if $s=4$, then $G_{1}^{\prime}$ has drop $(0.5,5)$ directly. In any case, $G_{1}^{\prime}$ has drop $(0.5,3^{*})$. We then simply need:
@@
Case III: $\boldsymbol{s\geq 2}$. By Proposition 39, we have $S(G_{2})\geq 1^{*}$; also, we must have $r\geq 5$, since if $r=4$ then $S(G_{0})\geq 2^{*}$ and it would covered in Case I. In particular, $\deg(x_{1})=4$.
We claim that $G-N[x_{1}]$ is not $3^{*}$-vul. For, suppose it were, and consider $u^{\prime}=x_{1}$ and its 3-neighbor $v_{1}^{\prime}=v_{1}$ and its two neighbors $x_{1}^{\prime}=u,x_{2}^{\prime}=x_{2}$. We have $r^{\prime}=\deg(x_{1}^{\prime})+\deg(x_{2}^{\prime})-2=4+\deg(x_{2})=r$. If $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])\leq 4-r^{\prime}$, then Proposition 40 applied to $u^{\prime}$ would give $R(N[x_{1}^{\prime}])=S(G_{0})\geq 2^{*}$ so $u$ would be covered in Case I. On the other hand, if $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])=\operatorname{minsurp}(G_{3}^{\prime})\geq 5-r^{\prime}$, then vertex $u^{\prime}$, with $v_{1}^{\prime},x_{1}^{\prime},x_{2}^{\prime}$, would be covered by Case II.
We claim next that $r=6$. For, suppose $r=5$, so $S(G_{0})\geq 1^{*}$. If $\operatorname{minsurp}(G_{3})\geq 0$, then we would be in Case II. On the other hand, if $\operatorname{minsurp}(G_{3})=-1$, then Proposition 40 gives $R(N[x_{1}])\geq 2^{*}$ and in particular $G-N[x_{1}]$ would be $3^{*}$-vul.
Now our branching rule is to generate subproblems $G_{0},G_{2},G_{3}$. By Proposition 40, we have $\operatorname{minsurp}(G_{3})\geq 0$ since $G-N[x_{1}]$ is not $3^{*}$-vul. So $G_{2}$ has drop $(3,8^{*})$ and we need:
@@
Case IV: $\boldsymbol{S(G_{0})\geq 3^{*}}$. Here, we generate subproblems $G_{0},G_{1}^{\prime}$. Since $s=1$, the graph $G_{1}^{\prime}$ has a subcubic vertex $v_{2}$ and so it has drop $(0.5,2^{*})$. We need:
@@
Case V: $\boldsymbol{r\geq 5}$ and $\boldsymbol{G_{0}}$ is $\boldsymbol{3^{*}}$-vul. Here we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$. We have $\operatorname{minsurp}(G_{3})\geq 5-r$, as otherwise Proposition 40 would give $R(N[x_{1}])\geq 3^{*}$ (bearing in mind that $s=1$); then $u^{\prime}=x_{1}$ would be covered in Case IV. So subproblem $G_{3}$ has drop $(2.5,7^{*})$. Overall, we need:
@@
Case VI: $\boldsymbol{r=6}$. Again we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$ as before. If $G-N[x_{1}]$ is $3^{*}$-vul, then vertex $u^{\prime}=x_{1}$ would be covered in Case V earlier (it has a 3-neighbor $v_{1}$ with a 4-neighbor $u$). So assume $G-N[x_{1}]$ is not $3^{*}$-vul; by Proposition 40, this implies $\operatorname{minsurp}(G_{3})\geq 0$ and so $G_{3}$ has drop $(3,8^{*})$. We need:
@@
Case VII: $\boldsymbol{r=4}$. Here, $G$ cannot have any 3-vertex $v$ with two 4-neighbors $y_{1},y_{2}$, as then vertex $u^{\prime}=y_{1}$ would have $r^{\prime}\geq 5$ and be covered in previous cases. So we assume that all 3-vertices in the graph have at least two 3-neighbors. In particular, we have $x_{1}\sim t_{1}$ and $x_{2}\sim t_{2}$ for 3-vertices $t_{1},t_{2}$ distinct from $v_{1}$. Then in $G_{1}^{\prime}$, the merged vertex $z_{1}$ has degree 4 and has 3-neighbors $t_{1},t_{2}$, that is, it is a special vertex (see Figure 6). Since $\operatorname{mindeg}(G_{1}^{\prime})\geq 3$ and $\operatorname{maxdeg}(G_{1}^{\prime})\leq 4$, we get $\operatorname{bonus}(G_{1}^{\prime})=\beta$. Also, $S(G_{0})\geq 2^{*}$ since $r=4$. We need:

### v7

@@ 6.1 Regularizing the graph and basic analysis
We will first ensure that $G$ has a number of nice properties:
@@
  • $G$ has no connected-components of size smaller than some constant $K$.
Let us explain the last clause here. Our branching algorithms will generate principal subproblems GXG-X with some given drop $(\Delta\mu,\Delta k)$ a constant $\Delta k$. If $K>5\Delta k+1$, then by Lemma 34, such subproblem automatically has drop $(\Delta\mu,\Delta k^{*})$ “for free”. We choose $K$ to be larger than all such constants needed for the analysis; we do not specify it explicitly.
@@
If $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 2$, but $G$ does not satisfy the Line-1 conditions, then it has available branch-seq $[(0,2)]$, $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $(1,3^{*}),(1,3^{*})]$
@@
If $G$ has a connected-component $X$ of size $O(1)$, we solve it exactly and update $k^{\prime}=k-|C_{X}|$ where $C_{X}$ is a good cover of $X$. Since $\operatorname{minsurp}(G)\geq 2$, we have $|C_{X}|\geq 2$ and this gives drop $[(0,2)]$.
If $G$ has a funnel $u$ with out-neighbor $x$, then we apply (B-Fun1), generating subproblems GN[u],GN[x]G-N[u],G-N[x]. Since $\operatorname{minsurp}(G)\geq 2$, these subproblems have drop $(1,3)$ directly; by Lemma 34 they indeed both have drop $(1,3^{*})$.
Otherwise, rules (P1) and (P2) cannot be applied; so for the remainder of the analysis we may suppose that $G$ is simplified.
@@
Finally, if $G$ has a 4-cycle $u_{1},u_{2},u_{3},u_{4}$, we consider two cases. First, suppose there is a “cross-edge”, say $u_{1}\sim u_{3}$. We must have $\deg(u_{1})=\deg(u_{3})=4$, else they would have a funnel. Let $v_{1},v_{3}$ be the other neighbors of $u_{1},u_{3}$. We split on $v_{1}$; subproblem $G-N[v_{1}]$ has drop $(1,3^{*})$ by Lemma 34. Subproblem $G-v_{1}$ has drop $(0.5,1)$ directly, and has a kite involving vertex $u_{1}$; by Lemma 14, after applying (P3), we get net drop $(1,3)$ and the resulting graph has the 3-vertex $u_{3}$.
Otherwise, suppose that $u_{1}\not\sim u_{3}$ and $u_{2}\not\sim u_{4}$. We apply (B-Shared) to get subproblems $G-\{u_{1},u_{3}\}$ and $G-\{u_{2},u_{4}\}$. These both have two non-adjacent subquadratic nodes ($u_{2},u_{4}$ and $u_{1},u_{3}$ respectively), so by Observation 12 they both have drop $(1,4)$. ∎
@@
After Line 1, for a 4-vertex $u$, the graph GN[u]G-N[u] has at least 8 subcubic vertices.
@@
Suppose $u$ has 3-neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{s}$ . We have $\operatorname{codeg}(u,v_{i})=0$ for $i\leq s$, since otherwise $v_{i}$ would have a 3-triangle. Likewise, we have $\operatorname{codeg}(u,v_{i})\leq 1$ for $i>s$, as otherwise $G$ would have a 4-cycle. Thus, the neighborhood of $u$ has at most $\lfloor(4-s)/2\rfloor$ edges, and hence at least $3s+4(4-s)-2\lfloor(4-s)/2\rfloor-4\geq 8$ edges from $N(u)$ to vertices $t$ at distance two from $u$. Since $G$ has no 4-cycles, each such vertex $t$ has an edge to precisely one vertex in $N(u)$. ∎
We now analyze the easier steps of the algorithm line-by-line. Observe that, starting at Line 4, the graph $G$ is 4-regular. Lines 3 and 5, which are more involved, will be considered later.
@@
First suppose that $\operatorname{bonus}(G)=0$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\geq 5$, then we apply Figure 6; otherwise, we apply Proposition 38. In order to satisfy Eq. (6), we need:
@@
Finally suppose $\operatorname{bonus}(G)=\alpha$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\leq 4$, then we can apply Proposition 38. Finally, suppose $G$ is simplified and has a vertex $u$ of degree at least $5$. If $\operatorname{shad}(N[u])\leq 4-\deg(u)$, then we apply Lemma 26. If $\operatorname{shad}(N[u])\geq 5-\deg(u)$, then we split on $u$; subproblem GN[u]G-N[u] has drop $(2,5)$, while subproblem GuG-u retains the subcubic vertex in the original graph and so has drop $(0.5,1^{*})$. Over all cases, we need:
@@
Line 4. Let $N(u)=\{v_{1},\dots,v_{4}\}$ where $v_{1}\sim v_{2}$. Since $G$ has no 4-cycles, $\{v_{1},v_{2}\}\not\sim\{v_{3},v_{4}\}$. We split on $v_{3}$, getting subproblems $G-N[v_{3}],G-v_{3}$ with drops $(1.5,4^{*})$ and $(0.5,1^{*})$. In the latter subproblem, we apply Proposition 30 to the funnel $u$ with out-neighbor $v_{4}$ and vertex $v_{1}\in N(u)\setminus N(v_{4})$, where $r=\deg_{G-v_{3}}(v_{1})+\deg_{G-v_{3}}(v_{4})-\operatorname{codeg}_{G-v_{3}}(v_{1},v_{4})-1\geq 4+3-1-1=5$. We need:
@@ 6.2 Line 3: Branching on a 4-vertex with a 3-neighbor
Suppose Line 3 selects a 4-vertex $u$, with 3 neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{4}$. We let $v=v_{1}$ be a chosen 3-vertex, and let $x,y$ be the other two neighbors of $v_{1}$, sorted so $\deg(x)\geq\deg(y)$. Since $G$ has no 4-cycles, and the 3-vertex $v$ has no triangles, all vertices $u,v_{i},x_{i},y_{i}$ are distinct, and $\{x,y\}\not\sim\{u,v_{2},v_{3},v_{4}\}$. See Figure 8.
Figure 8: The vertex $u$ and its associated neighbor vertices. The degrees of the nodes are shown.
Note that there may be multiple ways of ordering vertices in this way; e.g., we may permute $v_{1},\dots,v_{s}$, and if $\deg(x)=\deg(y)$ we can swap $x/y$. Let us fix some ordering. As a starting point, our strategy is to split on $u$, producing subproblems $G_{0}=G-N[u]$ and GuG-u. There are two main further branching strategies we can use on GuG-u:
  • Apply (P2), contracting its 2-vertices $v_{1},\dots,v_{s}$ into vertices $z_{1},\dots,z_{s}$; we denote the resulting graph by $G_{1}$. Overall we generate two subproblems $G_{0},G_{1}$.
  • Apply (B-Pair) to the 2-vertex $v$, generating further subproblems $G_{2}=G-N[v],G_{3}=G-u-N[x,y]$. Overall we generate three subproblems $G_{0},G_{2},G_{3}$.
Throughout Section 6.2 we write
$r$ $=\deg_{G_{1}}(z_{1})=\deg(x)+\deg(y)-2$
$s$ =# $=\text{\# of 3-neighbors of }u$
@@
For the first result, if $s\leq 3$, then $G-N[v_{1}]$ has a 3-vertex $v_{4}$ and a 2-vertex $v_{2}\not\sim v_{4}$; applying (P2) to $v_{2}$ reduces $k$ by one, and still leaves a subcubic vertex $v_{4}$. If $s=4$, then $G-N[v_{1}]$ has non-adjacent 2-vertices $v_{2},v_{3}$, so $S(G_{2})\geq 2$ by Proposition 10.
For the next two results, recall from Proposition 39 that $G_{0}$ has at least 8 subcubic vertices. If $r=4$, then $\deg_{G_{0}}(x)=\deg_{G_{0}}(y)=2$ and $x\not\sim y$, so by Proposition 33 we have $S(G_{0})\geq 2^{*}$. Similarly, if $r=5$, then $\deg_{G_{0}}(y)=2$ so by Proposition 32 we have $S(G_{0})\geq 1^{*}$. ∎
@@
If $r\geq 5$, then $G_{3}$ is a principal subproblem with excess two and $\Delta k=r+2$ and $\operatorname{minsurp}(G_{3})\geq 4-r$. Furthermore, the following bounds all hold:
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$, then $S(N[x])\geq 2^{*}$.
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$ and $s=1$, then $S(N[x])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-2$, then $S(N[x])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-1$, then GN[x]G-N[x] is $3^{*}$-vul.
@@
Here $\deg(x)=4$ and $\deg(y)=r-2$. Since $G$ has no 4-cycles, we have $|N[x,y]|=r+1$ and so $G_{3}$ is principal with excess two and $\Delta k=r+2$. If $s=1$, then the sole 3-neighbor of $u$ gets removed from $G_{3}$, and so $\operatorname{mindeg}(G_{3})\geq 1$. Also note that, by Proposition 39, the graph GN[x]G-N[x] has at least 8 subcubic vertices and $\operatorname{shad}(N[x])\geq 1$.
Let $I$ be a min-set in $G_{3}$ and $\ell=\operatorname{surp}_{G_{3}}(I)$. If $\ell\leq-2$ or $s=1$ then necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq\ell+(\deg_{G}(y)-1)=\ell+r-3$. We have $\operatorname{shad}(N[x])\geq 1$ so $\ell\geq 4-r$. Furthermore, if $\ell=4-r$ exactly, then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})=1$, and then Proposition 32 applied to indset $I\cup\{y\}$ gives $S(N[x])\geq 2^{*}$ for $|I|=1$ and $S(N[x])\geq 3^{*}$ for $|I|\geq 2$.
Finally, suppose $r=6$ and $\ell=-1$. If $|I|\geq 2$, then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 2$, which by Proposition 35 implies that GN[x]G-N[x] is $3^{*}$-vul. Otherwise, suppose $|I|=1$, i.e. $G_{3}$ has an isolated vertex $t$; since $G$ has no 4-cycles, this is only possible if $\deg_{G}(t)=3$ and $t\sim u$ and $\operatorname{codeg}(t,x)=\operatorname{codeg}(t,y)=1$. Then $\deg_{G-N[x]}(t)=2$ and by Proposition 32 we have $S(N[x])\geq 1^{*}$. ∎
@@
Case I: $\boldsymbol{s\geq 2}$ and $\boldsymbol{S(G_{0})\geq 2^{*}}$. We generate subproblems $G_{0},G_{1}$. If $s\in\{2,3\}$, then $G_{1}$ has drop $(0.5,3)$ and has a 3-vertex $v_{4}$; if $s=4$, then $G_{1}$ has drop $(0.5,5)$ directly. In any case, $G_{1}$ has drop $(0.5,3^{*})$. We then simply need:
@@
Case III: $\boldsymbol{s\geq 2}$. By Proposition 40, we have $S(G_{2})\geq 1^{*}$; also, we must have $r\geq 5$, since if $r=4$ then $S(G_{0})\geq 2^{*}$ and it would covered in Case I. In particular, $\deg(x)=4$.
We claim that GN[x]G-N[x] is not $3^{*}$-vul. For, suppose it were, and consider $u^{\prime}=x$ and its 3-neighbor $v^{\prime}=v$ and its two neighbors $x^{\prime}=u,y^{\prime}=y$. We have $r^{\prime}=\deg(x^{\prime})+\deg(y^{\prime})-2=r$. If $\operatorname{shad}(u^{\prime},N[x^{\prime},y^{\prime}])\leq 4-r^{\prime}$, then Proposition 41 applied to $u^{\prime}$ would give $S(N[x^{\prime}])=S(G_{0})\geq 2^{*}$ so $u$ would be covered in Case I. On the other hand, if $\operatorname{shad}(u^{\prime},N[x^{\prime},y^{\prime}])=\operatorname{minsurp}(G_{3}^{\prime})\geq 5-r^{\prime}$, then vertex $u^{\prime}$, with $v^{\prime},x^{\prime},y^{\prime}$, would be covered by Case II.
Now, if $\operatorname{minsurp}(G_{3})\leq-1$, then by Proposition 41, the graph GN[x]G-N[x] would be $3^{*}$-vul, which we have already ruled out. So $\operatorname{minsurp}(G_{3})\geq 0$. Consequently, if $r=5$ then observe that $S(G_{0})\geq 1^{*}$ and so we would be in Case II. So $r=6$, and our branching rule is to generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(3,8^{*})$ and we need:
@@
Case IV: $\boldsymbol{S(G_{0})\geq 3^{*}}$. We generate subproblems $G_{0},G_{1}$. Since $s=1$, the graph $G_{1}$ has a subcubic vertex $v_{2}$ and so it has drop $(0.5,2^{*})$. We need:
@@
Case V: $\boldsymbol{r\geq 5}$ and $\boldsymbol{G_{0}}$ is $\boldsymbol{3^{*}}$-vul. We generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$. We have $\operatorname{minsurp}(G_{3})\geq 5-r$, as otherwise Proposition 41 would give $S(N[x])\geq 3^{*}$ (bearing in mind that $s=1$); then $u^{\prime}=x$ would be covered in Case IV. So subproblem $G_{3}$ has drop $(2.5,7^{*})$. Overall, we need:
@@
Case VI: $\boldsymbol{r\geq 5}$. We must have specifically $r=6$, as if $r=5$ then by Proposition 41 we would have $S(G_{0})\geq 1^{*}$ putting us into Case V.
Again we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$ as before. If GN[x]G-N[x] is $3^{*}$-vul, then vertex $u^{\prime}=x$ would be covered in Case V earlier. So assume GN[x]G-N[x] is not $3^{*}$-vul; by Proposition 41, this implies $\operatorname{minsurp}(G_{3})\geq 0$ and so $G_{3}$ has drop $(3,8^{*})$. We need:
@@
Case VII: Everything else. Here, $G$ cannot have any 3-vertex with two 4-neighbors, as any such 4-neighbor $u^{\prime}$ would have $r^{\prime}\geq 5$ and be covered in Case VI. So we assume that all 3-vertices in the graph have at least two 3-neighbors. In particular $r=4$, and $x\sim x^{\prime}$ and $y\sim y^{\prime}$ for 3-vertices $x^{\prime},y^{\prime}$ distinct from $v$. We must have $x^{\prime}\neq y^{\prime}$ since $G$ has no 4-cycles. The merged vertex $z_{1}$ in $G_{1}$ has degree 4 with 3-neighbors $x^{\prime},y^{\prime}$, that is, it is a special vertex (see Figure 9). Since $\operatorname{mindeg}(G_{1})\geq 3$ and $\operatorname{maxdeg}(G_{1})\leq 4$, we get $\operatorname{bonus}(G_{1})=\beta$. Also, $S(G_{0})\geq 2^{*}$ since $r=4$. We need: